# The amplitude decay of a harmonic oscillator damped simultaneously by weak linear and nonlinear damping forces

### v1

@@ V Discussion and conclusion
By comparing our analytical solutions with numerical solutions, we have shown that our analytical solutions give an excellent description of the dynamics of a harmonic oscillator simultaneously damped by sliding friction and damping forces that are linear and quadratic in velocity if $c_{0}+c_{1}+c_{2}\ll\omega_{0}$ is fulfilled. As a paradigmatic example, we considered the block-spring system with damping force (1). An example of an experiment with the block-spring system damped (only) by sliding friction is given in Kamela , where oscillations of a system consisting of two carts and a wooden block on a track were considered. In such an experiment, sliding friction corresponding to small values of $c_{0}$ (e.g. to $c_{0}\sim 0.01\omega_{0}$) could be easily realized by taking a lighter wooden block or completely removing it and leaving only the cart to oscillate, since the damping of the cart itself corresponds to weak sliding friction Kamela . By adding, for example, small sails to such a system, to increase air resistance, a combination of all three damping forces can be realized, and the values of the coefficients $c_{1}$ and $c_{2}$ can be influenced by choosing the shape and size of the sail Wang , as well as the initial conditions, since $c_{2}$ depends on $A_{0}$. Our results can, of course, be easily applied to all other physical systems described by the damped harmonic oscillator equation. For example, in Wang all three types of damping were individually experimentally investigated with physical pendulums. The harmonic oscillator damped by constant damping was realized by a physical pendulum with sliding friction in the support, oscillations with linear damping were realized by magnetically damped physical pendulum (damping by eddy currents), and oscillations of a physical pendulum with (dominantly) quadratic damping were achieved with air resistance Wang . The values of the physical parameters and initial conditions used in Wang correspond to $c_{0}\ll\omega_{0}$, $c_{1}\ll\omega_{0}$, and $c_{2}\ll\omega_{0}$. Therefore, using a physical pendulum, with a combination of the mentioned effects, it is possible to design a system that corresponds to a harmonic oscillator simultaneously damped by three damping forces and with $c_{0}+c_{1}+c_{2}\ll\omega_{0}$ fulfilled. Thus, our theoretical results can be tested experimentally.
@@
In conclusion, we showed how to obtain an excellent analytical approximation of the amplitude decay, i.e. of the solutions that describe the dynamics of harmonic oscillator simultaneously damped with linear and nonlinear damping forces. Our results are suitable for undergraduate physics classes, since we use simple mathematics that is well known to undergraduate students. Based on our theoretical results, laboratory exercises can be designed in which an interesting task for students could be the measurement of the duration of free oscillations and the analysis of the validity of approximate analytical expressions (20)-(22) depending on the parameters of the system and initial conditions. Furthermore, vibrations simultaneously damped with linear and nonlinear damping forces are a topic of wider interest that goes beyond physics teaching. For example, a paper JSVCoulombViscous was recently published in which oscillations with a combination of sliding friction and linear damping were considered due to their importance in the context of engineering applications and the amplitude decay of the same form as the function (28) was derived. Since the approach we use is mathematically simpler than the phase space coordinate averaging used in JSVCoulombViscous , and also easily enables the addition of a third damping force, i.e. quadratic damping, we believe that it can be useful in the context of applications in engineering as well.
VI Acknowledgments

### v2

@@ VI Discussion and conclusion
By comparing our analytical solutions with numerical solutions, we have shown that our analytical solutions give an excellent description of the dynamics of a harmonic oscillator simultaneously damped by sliding friction and damping forces that are linear and quadratic in velocity if $c_{0}+c_{1}+c_{2}\ll\omega_{0}$ is fulfilled. As a paradigmatic example, we considered the block-spring system with damping force (1). An example of an experiment with the block-spring system damped (only) by sliding friction is given in Kamela , where oscillations of a system consisting of two carts and a wooden block on a track were considered. In such an experiment, sliding friction corresponding to small values of $c_{0}$ (e.g. to $c_{0}\sim 0.01\omega_{0}$) could be easily realized by taking a lighter wooden block or completely removing it and leaving only the cart to oscillate, since the damping of the cart itself corresponds to weak sliding friction Kamela . By adding, for example, small sails to such a system, to increase air resistance, a combination of all three damping forces can be realized, and the values of the coefficients $c_{1}$ and $c_{2}$ can be influenced by choosing the shape and size of the sail Wang , as well as the initial conditions, since $c_{2}$ depends on $A_{0}$. Our results can, of course, be easily applied to all other physical systems described by the damped harmonic oscillator equation. For example, in Wang all three types of damping were individually experimentally investigated with physical pendulums. The harmonic oscillator damped by constant damping was realized by a physical pendulum with sliding friction in the support, oscillations with linear damping were realized by magnetically damped physical pendulum (damping by eddy currents), and oscillations of a physical pendulum with (dominantly) quadratic damping were achieved with air resistance Wang . The values of the physical parameters and initial conditions used in Wang correspond to $c_{0}\ll\omega_{0}$, $c_{1}\ll\omega_{0}$, and $c_{2}\ll\omega_{0}$. Therefore, using a physical pendulum, with a combination of the mentioned effects, it is possible to design a system that corresponds to a harmonic oscillator simultaneously damped by three damping forces and with $c_{0}+c_{1}+c_{2}\ll\omega_{0}$ fulfilled. Thus, our theoretical results can be tested through experiments suitable for students at the undergraduate level.
@@
In conclusion, we showed how to obtain an excellent analytical approximation of the amplitude decay, i.e. of the solutions that describe the dynamics of harmonic oscillator simultaneously damped with linear and nonlinear damping forces. Our results are suitable for undergraduate physics courses, since we use basic mathematical concepts familiar to students at this level. The presented approach is a good example of the importance of modeling in physics, i.e. it can be emphasized to students that even when strong assumptions are used, models can be obtained that give a very good quantitative description of complex phenomena. Based on our theoretical results, laboratory exercises can be designed in which an interesting task for students could be the measurement of the duration of free oscillations and the analysis of the validity of approximate analytical expressions (21)-(23) depending on the parameters of the system and initial conditions. Furthermore, vibrations simultaneously damped with linear and nonlinear damping forces are a topic of wider interest that goes beyond physics teaching. For example, a paper JSVCoulombViscous was recently published in which oscillations with a combination of sliding friction and linear damping were considered due to their importance in the context of engineering applications and the amplitude decay of the same form as the function (29) was derived. Since the approach we use is mathematically simpler than the phase space coordinate averaging used in JSVCoulombViscous , and also easily enables the addition of a third damping force, i.e. quadratic damping, we believe that it can be useful in the context of applications in engineering as well. Finally, in the last paragraph of Section V, we commented on possible improvements to the approach presented in this paper, which will be the topic of our future work.
VII Acknowledgments